# RotBench: Evaluating Multimodal Large Language Models on Identifying Image Rotation

### v1

@@ Abstract
Tianyi Niu  Jaemin Cho  Elias Stengel-Eskin  Mohit Bansal UNC Chapel Hill
@@ 2 Related Work
Sensitivity to visual perturbations.  Previous work has shown vision encoders and MLLMs are sensitive to simple image transformations. Anis et al. (2025) evaluate CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023) on a suite of common image transformations – rotations, flips, noise, etc. – revealing substantial gaps between human and model understanding. Usama et al. (2025) finds MLLMs exhibit distinct failure patterns in scene-text and object reasoning tasks when applying ImageNet-C corruptions Hendrycks and Dietterich (2019) to image inputs. These studies highlight that despite strong clean-image performance, vision encoders and MLLMs are highly sensitive to photometric and geometric distortions.
@@
Image orientation estimation.  Fine-tuning models to identify image orientation has been the focus of prior work Xu et al. (2024). For example, Fischer et al. (2015) and Joshi and Guerzhoy (2017) focused on fine-tuning convolutional neural networks (CNNs) to estimate and identify image rotation. While our work tackles a similar task, we are instead interested in the problem as a test of general-purpose MLLMs’ inherent reasoning abilities, i.e., whether they can estimate image rotation without extensive fine-tuning. Note that our work aligns more closely with the problem of image, not camera, orientation estimation.
Camera orientation estimation  Camera orientation estimation is also a well-studied task in computer vision Xian et al. (2019). Instead of predicting the rotation of the image, camera orientation estimation seeks to predict the spatial location of the camera when capturing an image. Contemporary approaches of this task use deep networks to directly predict orientation parameters from image features in an end-to-end manner Xian et al. (2019); Lee et al. (2021, 2020).
@@ 4.3 Auxiliary Information
Figure˜8 illustrates all forms of auxiliary information provided to the model. Note that all auxiliary information is separately extracted for each rotation, ensuring our approach does not depend on prior knowledge of the image’s orientation.
@@ 5 Main Results
MLLMs accurately identify right-side-up (0°) images.  All evaluated models effectively recognize right-side-up (0°) images. Qwen-2.5-VL-7B-Instruct (Qwen) achieves an accuracy of 0.99 without supplemental data. Proprietary models (GPT-4o, GPT-4.1, o3, Gemini-2.5-Flash, Gemini-2.5-Pro, and Gemini-2.0-Flash) consistently exhibit near-perfect accuracy on identifying unrotated images. This outcome aligns with expectations, given these models likely encountered predominantly upright images during training, and thus 0° can be assumed to be the default option.
@@
Proprietary models perform well on upside-down (180°) images.  All models except Qwen demonstrate robust performance on images rotated 180°, with GPT-4o, GPT-4.1, o3, and Gemini-2.5-Pro all achieving accuracies notably above chance (> 0.7). However, Gemini-2.0-Flash and Gemini-2.5-Flash display relatively lower zero-shot performance, with accuracies around 0.5. This indicates that state-of-the-art proprietary models generally possess a reliable capability to recognize upside-down images, though there remains variability within the different model families.
@@
Identifying 90° and 270° is challenging for all evaluated models.  All models exhibit substantial difficulties when distinguishing between 90° and 270°, while the poorest performance consistently emerged with 270° images. Confusion matrix analysis (Section˜6.1) reveals frequent misclassifications between these two orientations, indicating a distinct challenge when identifying 0° and 180°.
@@
Providing auxiliary inputs does not reliably improve performance.  None of the auxiliary information results in meaningful, consistent performance gains across all tested models. Paradoxically, the introduction of additional information sometimes leads to marginal performance degradation. When including all forms of auxiliary information, Qwen’s accuracy on 90° decreased from 0.51 to 0.26. Similarly, Gemini-2.0-Flash’s accuracy on 270° decreased from 0.44 to 0.17.
@@
Rotation grid improves performance only for reasoning models.  Providing the rotation grid degraded performance on most models. Gemini-2.0-Flash accuracy on 90° decreased by nearly 0.5 compared to CoT prompting. Rotation grid guided, however, improved GPT-4o accuracy on 270° by around 0.1. At the same time, both o3 and Gemini-2.5-Pro saw performance improvements. Notably, Gemini-2.5-Pro’s accuracy on 90° and 270° improved by 0.15. These results suggest the two reasoning models are much more effective at utilizing visual context. We further leverage the robust identification of 0° images and modify the rotation grid into a majority voting approach Appendix˜I. This setup results in gains even in weaker models, indicating MLLMs can achieve moderate performance with sufficient visual scaffolding.
@@
Chain-of-thought improves 180° performance, but results are mixed on 90° and 270°.  Employing chain-of-thought (CoT) prompting yields mixed results, differing notably across models and orientations. Gemini models show improved accuracy on 90° rotations but decreased accuracy for 270°. Conversely, GPT models exhibit improved performance on 270° rotations, coupled with degraded accuracy for 90° rotations. However, CoT consistently enhanced accuracy for 180° rotations across all models tested. These findings suggest CoT prompting can help models better reason about rotations to some extent, but does not universally resolve inherent challenges in distinguishing portrait orientations. Appendix˜F examines a common failure of CoT in detail.
@@ 6.1 Model Bias Towards 0° and 90°
To further elucidate the specific types of rotational errors made by models, we analyze the confusion matrix for GPT-4o. Figure˜3 shows the confusion matrix obtained from summing predictions across three runs. We see GPT-4o predominantly struggles with differentiating between 90° and 270° rotations. Specifically, the model misclassifies 459 instances of 90° images as 270° and 424 instances of 270° images as 90°. Yet, 0° and 180° show significantly fewer misclassifications. This analysis underscores a critical shortcoming in model performance, suggesting either that the vision encoder is not providing sufficient signals to distinguish between clockwise and counter-clockwise rotations, or that the MLLM is not adequately incorporating the visual information into its reasoning.
@@ 6.2 Distinguishing Clockwise from Counter-clockwise Rotations
Following the analysis in Section˜6.1, we specifically examine whether MLLMs can distinguish clockwise (CW) and counter-clockwise rotations (CCW). For this, we simplified the original four-way classification task into a binary classification task. Using the 90° and 270° images from RotBench-large, we ask MLLMs to determine whether an image has been rotated 90° CCW or CW. [^6] Note that 90° CW is equivalent to a 270° CCW. By asking models to explicitly differentiate the two directions of rotation, we hope to provide a clearer signal for directional understanding. Figure˜4 shows an example question where the GPT-4o provides an incorrect answer.
@@
These findings strongly indicate that the MLLMs tested have limitations in reliably distinguishing between CW and CCW rotational directions, offering a potential explanation for why distinguishing 90° and 270° CCW rotations presents such a challenging task. Further corroborating this idea, we show that evaluating Qwen-2.5-VL-7B-Instruct, GPT-4o, o3 on RotBench-small using clockwise angles does not result in significant performance differences (Appendix˜D). Furthermore, in Appendix˜H we find that even with in-context examples, models struggle to correctly discriminate between clockwise and counterclockwise-rotated images.
@@ 6.3 Can Fine-tuning Solve RotBench?
Figure 5: Qwen-2.5-VL-7B-Instruct accuracy on different degrees of rotation as training progresses.
@@
Results.Fig.˜5 reveals a high and consistent accuracy for 0° throughout training, indicating robust recognition of upright images. Performance on 180° gradually improves, stabilizing around 0.8 after approximately 7000 images. However, accuracies for 90° and 270° exhibit substantial oscillations, suggesting that the model did not achieve stable improvement. The model appears to be caught in a cycle, alternating between accuracy gains and losses for these two rotations. This phenomenon is also reflected in our main results (Section˜5), where using CoT prompting improved accuracy on 270° images at the expense of 90°. The unstable performance may indicate potential representational constraints in current visual encoders that limit visual understanding capabilities, particularly regarding subtle rotational distinctions.
@@ A.1 Further details on Stage 1 and 2 Filtering
Stage 1.  We randomly sample 300 images from Spatial-MM Shiri et al. (2024). Spatial-MM is divided into two splits, one-subject and two-subject, respectively comprised of images that contain one and two primary subjects. As models may use image dimensions to infer orientation, we crop each image into a square before further processing. We sample 150 images from each split and show the resulting 300-image dataset to the Stage 1 annotator. For each image, the annotator can decide to either accept, discard, or flag. The annotator accepts an image if it (1) contains easily identifiable rotational visual cues (e.g., a person standing), and (2) the image has meaningful differences when rotated. Images that do not satisfy these criteria are discarded. Occasionally, some images have subtle visual cues, or require more detailed semantic understanding to correctly perceive orientation. These images tend to have primary subjects that do not vary significantly with rotation, or require integrating background signals to identify rotation. The sample image shown in Fig.˜6 is an example of a flagged image. Such images are flagged by the Stage 1 annotator to proceed to Stage 2. We provide further examples and details of accepted, flagged, and discarded images in Fig.˜9. Note that the Stage 1 annotator is also tasked with ensuring no leaked personal information or offensive content is incorporated into RotBench.
@@ A.4 Annotator Interface
Figure˜7 shows the interface provided to Stage 2 annotators. The mapping between letter choice and rotation degree is shuffled from question to question.
@@ A.5 Sample Images
Figure˜8 displays the various forms of additional information provided to the model. Figure˜9 provides examples of an image that is accepted, flagged, and discarded during Stage 1 of data filtering. The accepted image has a clear subject and is easily distinguishable when rotated 0°, 90°, 180°, and 270°. The flagged image is more difficult to identify, however, the slight tilt – as opposed to a directly top-down perspective – still enables accurate judgment of rotation. However, the discarded image has no meaningful signals to distinguish between the four orientations.
@@ A.6 Further details on flagged images
Flagged images typically involve cases where there is a lack of primary subjects or the primary subject does not offer any distinguishable signals under rotation. These images tend to require deeper comprehension in order to make an accurate classification. For instance, Fig.˜6 displays a blank projector canvas above some colorful couches on a patio. The orientation of a projector screen cannot be reliably identified by where its supporting structures are located – the screen can be hung (cable hangs down), pitched up (supporting stand on the ground), or attached to a wall (support frame located sideways of the screen). Therefore, correctly identifying the orientation of the image requires understanding the correct orientation of the couches.

### v2

@@ Abstract
Tianyi Niu  Jaemin Cho  Elias Stengel-Eskin  Mohit Bansal UNC Chapel Hill {tianyin4, jmincho, esteng, mbansal}@cs.unc.edu
@@ 2 Related Work
Sensitivity to visual perturbations.  Previous work has shown that visual encoders and MLLMs are sensitive to simple image transformations. Anis et al. (2025) evaluate CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023) on a suite of common image transformations – rotations, flips, noise, etc. – revealing substantial gaps between human and model understanding. Usama et al. (2025) finds MLLMs exhibit distinct failure patterns in scene-text and object reasoning tasks when applying ImageNet-C corruptions Hendrycks and Dietterich (2019) to image inputs. Taken together, these studies highlight that despite strong clean-image performance, visual encoders and MLLMs are highly sensitive to photometric and geometric distortions.
@@
Identifying visual perturbations.  Past work has tested how well vision encoders and MLLMs can identify which perturbation or transformation was applied to an input image (Lin et al., 2020; An et al., 2023; Rashtchian et al., 2023). Udandarao et al. (2024) show that visual encoders and MLLMs often fail to classify which perturbation out of 27 perturbations (e.g., contrast, brightness, rotation, blur) was applied to an input image on automatically-created datasets. In contrast, we provide a fine-grained analysis of rotation specifically through a 4-way angle classification task on a curated dataset with rigorous human verification, offering novel insights into why state-of-the-art modern MLLMs struggle with distinguishing specific rotation angles. Moreover, we test providing remedies such as chain-of-thought reasoning, various forms of auxiliary information (e.g., captions, bounding boxes, segmentation maps, depth maps), and guiding reasoning with algebraic calculations. We also analyze whether model performance is dependent on abilities such as identifying clockwise versus counter-clockwise rotation.
Image and camera orientation estimation.  Fine-tuning models to identify image orientation has been the focus of prior work Xu et al. (2024). For example, Fischer et al. (2015) and Joshi and Guerzhoy (2017) focused on fine-tuning convolutional neural networks (CNNs) to estimate and identify image rotation. While our work tackles a similar task, we are instead interested in the problem as a test of general-purpose MLLMs’ inherent reasoning abilities, i.e., whether they can estimate image rotation without extensive fine-tuning. Note that our work aligns more closely with the problem of image, not camera, orientation estimation. Camera orientation estimation is also a well-studied task in computer vision Xian et al. (2019). Instead of predicting the rotation of the image, camera orientation estimation seeks to predict the spatial location of the camera when capturing an image. Contemporary approaches of this task use deep networks to directly predict orientation parameters from image features in an end-to-end manner Xian et al. (2019); Lee et al. (2021, 2020).
@@ 4.3 Auxiliary Information
Figure˜3 illustrates all forms of auxiliary information provided to the model. Note that all auxiliary information is separately extracted for each rotation, ensuring our approach does not depend on prior knowledge of the image’s orientation.
@@ 5 Main Results
MLLMs accurately identify right-side-up (0°) images.
@@
All evaluated models effectively recognize right-side-up (0°) images. Qwen-2.5-VL-7B-Instruct (Qwen) achieves an accuracy of 0.99 without supplemental data. Proprietary models (GPT-4o, GPT-4.1, o3, Gemini-2.5-Flash, Gemini-2.5-Pro, and Gemini-2.0-Flash) consistently exhibit near-perfect accuracy on identifying unrotated images. This outcome aligns with expectations, given these models likely encountered predominantly upright images during training, and thus 0° can be assumed to be the default option.
Proprietary models perform well on upside-down (180°) images.
@@
All models except Qwen demonstrate robust performance on images rotated 180°, with GPT-4o, GPT-4.1, o3, and Gemini-2.5-Pro all achieving accuracies notably above chance (> 0.7). However, Gemini-2.0-Flash and Gemini-2.5-Flash display relatively lower zero-shot performance, with accuracies around 0.5. This indicates that state-of-the-art proprietary models generally possess a reliable capability to recognize upside-down images, though there remains variability within the different model families.
Identifying 90° and 270° is challenging for all evaluated models.
@@
All models exhibit substantial difficulties when distinguishing between 90° and 270°, while the poorest performance consistently emerged with 270° images. Confusion matrix analysis (Section˜6.1) reveals frequent misclassifications between these two orientations, indicating a distinct challenge when identifying 0° and 180°.
Providing auxiliary inputs does not reliably improve performance.
@@
None of the auxiliary information results in meaningful, consistent performance gains across all tested models. Paradoxically, the introduction of additional information sometimes leads to marginal performance degradation. When including all forms of auxiliary information, Qwen’s accuracy on 90° decreased from 0.51 to 0.26. Similarly, Gemini-2.0-Flash’s accuracy on 270° decreased from 0.44 to 0.17.
Rotation grid improves performance only for reasoning models.
@@
Providing the rotation grid degraded performance on most models. Gemini-2.0-Flash accuracy on 90° decreased by nearly 0.5 compared to CoT prompting. Rotation grid guided, however, improved GPT-4o accuracy on 270° by around 0.1. At the same time, both o3 and Gemini-2.5-Pro saw performance improvements. Notably, Gemini-2.5-Pro’s accuracy on 90° and 270° improved by 0.15. These results suggest the two reasoning models are much more effective at utilizing visual context. We further leverage the robust identification of 0° images and modify the rotation grid into a majority voting approach Appendix˜I. This setup results in gains even in weaker models, indicating MLLMs can achieve moderate performance with sufficient visual scaffolding.
CoT improves 180° performance, but results are mixed on 90° and 270°.
@@
Employing chain-of-thought (CoT) prompting yields mixed results, differing notably across models and orientations. Gemini models show improved accuracy on 90° rotations but decreased accuracy for 270°. Conversely, GPT models exhibit improved performance on 270° rotations, coupled with degraded accuracy for 90° rotations. However, CoT consistently enhanced accuracy for 180° rotations across all models tested. These findings suggest CoT prompting can help models better reason about rotations to some extent, but does not universally resolve inherent challenges in distinguishing portrait orientations. Appendix˜F examines a common failure of CoT in detail.
@@ 6.1 Model Bias Towards 0° and 90°
To further elucidate the specific types of rotational errors made by models, we analyze the confusion matrix for GPT-4o. Figure˜4 shows the confusion matrix obtained from summing predictions across three runs. We see GPT-4o predominantly struggles with differentiating between 90° and 270° rotations. Specifically, the model misclassifies 459 instances of 90° images as 270° and 424 instances of 270° images as 90°. Yet, 0° and 180° show significantly fewer misclassifications. This analysis underscores a critical shortcoming in model performance, suggesting either that the vision encoder is not providing sufficient signals to distinguish between clockwise and counter-clockwise rotations, or that the MLLM is not adequately incorporating the visual information into its reasoning.
@@ 6.2 Distinguishing Clockwise from Counter-clockwise Rotations
Following the analysis in Section˜6.1, we specifically examine whether MLLMs can distinguish clockwise (CW) and counter-clockwise rotations (CCW). For this, we simplified the original four-way classification task into a binary classification task. Using the 90° and 270° images from RotBench-large, we ask MLLMs to determine whether an image has been rotated 90° CCW or CW. [^6] Note that 90° CW is equivalent to a 270° CCW. By asking models to explicitly differentiate the two directions of rotation, we hope to provide a clearer signal for directional understanding. Figure˜5 shows an example question where the GPT-4o provides an incorrect answer.
@@
These findings strongly indicate that the MLLMs tested have limitations in reliably distinguishing between CW and CCW rotational directions, offering a potential explanation for why distinguishing 90° and 270° CCW rotations presents such a challenging task. Further corroborating this idea, we show that evaluating Qwen-2.5-VL-7B-Instruct, GPT-4o, o3 on RotBench-small using clockwise angles does not result in significant performance differences (Appendix˜D).
@@ 6.3 Can Fine-tuning Solve RotBench?
Figure 6: Qwen-2.5-VL-7B-Instruct’s accuracy on different degrees of rotation as training progresses.
@@
Results.Fig.˜6 reveals a high and consistent accuracy for 0° throughout training, indicating robust recognition of upright images. Performance on 180° gradually improves, stabilizing around 0.8 after approximately 7000 images. However, accuracies for 90° and 270° exhibit substantial oscillations, suggesting that the model did not achieve stable improvement. The model appears to be caught in a cycle, alternating between accuracy gains and losses for these two rotations. This phenomenon is also reflected in our main results (Section˜5), where using CoT prompting improved accuracy on 270° images at the expense of 90°. The unstable performance may result from potential representational constraints in current visual encoders that limit visual understanding capabilities, particularly regarding subtle rotational distinctions.
@@ A.1 Further details on Stage 1 and 2 Filtering
Stage 1.  We randomly sample 300 images from Spatial-MM Shiri et al. (2024). Spatial-MM is divided into two splits, one-subject and two-subject, respectively comprised of images that contain one and two primary subjects. As models may use image dimensions to infer orientation, we crop each image into a square before further processing. We sample 150 images from each split and show the resulting 300-image dataset to the Stage 1 annotator. For each image, the annotator can decide to either accept, discard, or flag. The annotator accepts an image if it (1) contains easily identifiable rotational visual cues (e.g., a person standing), and (2) the image has meaningful differences when rotated. Images that do not satisfy these criteria are discarded. Occasionally, some images have subtle visual cues, or require more detailed semantic understanding to correctly perceive orientation. These images tend to have primary subjects that do not vary significantly with rotation, or require integrating background signals to identify rotation. The sample image shown in Fig.˜7 is an example of a flagged image. Such images are flagged by the Stage 1 annotator to proceed to Stage 2. We provide further examples and details of accepted, flagged, and discarded images in Fig.˜9. Note that the Stage 1 annotator is also tasked with ensuring no leaked personal information or offensive content is incorporated into RotBench.
@@ A.4 Annotator Interface
Figure˜8 shows the interface provided to Stage 2 annotators. The mapping between letter choice and rotation degree is shuffled from question to question.
@@ A.5 Sample Images
Figure˜3 displays the various forms of additional information provided to the model. Figure˜9 provides examples of an image that is accepted, flagged, and discarded during Stage 1 of data filtering. The accepted image has a clear subject and is easily distinguishable when rotated 0°, 90°, 180°, and 270°. The flagged image is more difficult to identify, however, the slight tilt – as opposed to a directly top-down perspective – still enables accurate judgment of rotation. However, the discarded image has no meaningful signals to distinguish between the four orientations.
@@ A.6 Further details on flagged images
Flagged images typically involve cases where there is a lack of primary subjects or the primary subject does not offer any distinguishable signals under rotation. These images tend to require deeper comprehension in order to make an accurate classification. For instance, Fig.˜7 displays a blank projector canvas above some colorful couches on a patio. The orientation of a projector screen cannot be reliably identified by where its supporting structures are located – the screen can be hung (cable hangs down), pitched up (supporting stand on the ground), or attached to a wall (support frame located sideways of the screen). Therefore, correctly identifying the orientation of the image requires understanding the correct orientation of the couches.